Write the augmented matrix for the system of linear equations
3x-2y+5z=31
x+3y-3z=-12
-2x-5y+3z=11
Thanks in advance

Answer by jim_thompson5910(35256)

The system of linear equations

3x-2y+5z=31
x+3y-3z=-12
-2x-5y+3z=11

converts to augmented matrix